OpenStudy (anonymous):

What is the volume of a cube that has sides 9.2 inches long?

OpenStudy (danjs):

The cube of that length

OpenStudy (usukidoll):

Volume of a cube formula is Volume = s^3 where s = side so Volume=(9.2)^3 or Volume = 9.2 x 9.2 x 9.2

OpenStudy (anonymous):

you will use this equation to find the volume of cube \[x^{3}\] so we have x_9.2 so it will lead us to \[9.2^{3}\]